OpenStudy (anonymous):

The plane x + y + 2z = 10 intersects the paraboloid z = x^2 + y^2 in an ellipse. Find the points on this ellipse that are nearest to and farthest from the origin.

OpenStudy (anonymous):

gradient f = <2x, 2y, 2z> gradient g = <1,1,2> gradient h = <2x, 2y, -1> 2x=λ+2vx 2y=λ+2vy 2z=2λ-v x+y+2z=10 \[x^2+y^2-z=0\] 2x-2xv=λ 2y-2yv=λ (2z+v)/2=λ 2x-2xv=2y-2yv 2x(1-v)=2y(1-v) either x=y or v=1 case 1: x=y x+x+2z=16 == 2x+2z=16 x^2+x^2-z=0 == 2x^2-x=0 2x=16-2z x=8-z 2(8-z)^2-z=0 2(z^2-16z+64)-z=0 2z^2-32z+128-z=0 2z^2-33z+128=0 not sure where to go from here or what i did wrong case 2: v =1 2x=λ+2x λ=0 2z=2(0)-(1) 2z = -1 z = -1/2 x+y+2(-1/2) = 10 x+y=9 x=9-y x^2+y^2-(-1/2)=0 (9-y)^2+y^2+1/2=0 y^2-1881+y^2+1/2=0 2y^2-18y+81+1/2=0 and once again im stuck here as well

OpenStudy (anonymous):